What are the two factors of -63 that add up to -2?
Answer by Boreal(15235)

xy=-63
x+y=-2
x=-y-2
substitute
(-y-2)y=-63
-y^2-2y=63
0=y^2+2y-63
0=(y-7)(y+9)
y=7,-9
They add to -2 and multiply to -63